OpenStudy (anonymous):

I need help proving this trigonometric identity step-by-step? (Will give medal and fan) tan^2 x - sin^2 x = tan^2 x sin^2 x I'm given two columns: Calculation and Reason. I was given the reasons (which I will list in order in the comments), but I need to list my calculations according to the reasons.

OpenStudy (anonymous):

Reasons: 1. This follows the definition of tan x = sin x / cos x 2. This follows from factoring out sin^2 x 3. This follows by the definition of sec x = 1/cos x 4. This follows from the Pythagorean identity sec^2 x - 1 = tan^2 x

OpenStudy (anonymous):

@whpalmer4 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

oops last step must be (sin^2x * sin^2x) / cos^2x

OpenStudy (anonymous):

then we will get tan^2x *sin^2x

OpenStudy (anonymous):

Have you understood ?

OpenStudy (anonymous):

Can you explain to me how those steps refer to the reasons? Like which step corresponds with which reason? Sorry or the late response

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

At left hand side we replace tan^2x by sec^2x - 1 okay ?

OpenStudy (anonymous):

But the first reason says I'm supposed to use the identity tanx = sinx/cosx, does that go along with that reason?

OpenStudy (anonymous):

this is trigonometric formula ..

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

Okay we are solve using this tanx = sinx/cosx So tan^2x = sin^2x / cos^x So replace tan^2x by sin^2x / cos^2x at left hand side ,, what do you get ?

OpenStudy (anonymous):

sin^2 x / cos^2 x - sin^2 x ?

OpenStudy (anonymous):

correct !! now take LCM

OpenStudy (anonymous):

what's LCM?

hero (hero):

The truth is, different people have different methods for figuring out trigonmetric identities. In this case, while some would choose to take the route of converting to \(\sec^2x\) to finish solving, I am perfectly fine with producing the following set of steps: \(\dfrac{\sin^2x}{\cos^2x} - \sin^2x\) Factor out \(\sin^2x\) to get: \(\sin^2x\left(\dfrac{1}{\cos^2x} - 1\right)\) Re-write 1 as \(\dfrac{\cos^2x}{\cos^2x}\): \(\sin^2x\left(\dfrac{1}{\cos^2x} - \dfrac{\cos^2x}{\cos^2x}\right)\) Combine fractions: \(\sin^2x\left(\dfrac{1-\cos^2x}{\cos^2x}\right)\) Replace \(1 - \cos^2x\) with \(\sin^2x\): \(\sin^2x\left(\dfrac{\sin^2x}{\cos^2x}\right)\) \(\sin^2x \tan^2x\)

OpenStudy (anonymous):

Is that the finished proof? Thank you so much!

hero (hero):

No, that still isn't what you're looking for, but I can demonstrate how they wanted you to do it. I was just expressing a point that they are forcing you to do it "only one way" which I don't exactly agree with.

hero (hero):

What they wanted you to do is this: \(\dfrac{\sin^2x}{\cos^2x} - \sin^2x\) Factor out \(\sin^2x\): \(\sin^2x\left(\dfrac{1}{\cos^2x} - 1\right)\) Observe that \(\dfrac{1}{\cos^2x} = \dfrac{1}{\cos(x)}\dot\ \dfrac{1}{\cos(x)}\) \(\sin^2x\left(\dfrac{1}{\cos(x)}\dot\ \dfrac{1}{\cos(x)} - 1\right)\) Recall that \(\dfrac{1}{\cos(x)} = \sec(x)\): \(\sin^2x\left(\sec(x)\sec(x)- 1\right)\) \(\sin^2x\left(\sec^2x - 1\right)\) And finally by Pythagorean Identity: \(\sin^2x\tan^2x\)

hero (hero):

Something like that